• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 19:32:50 on February 06, 2009

    Entry number 259321

    This entry is a followup to: 259133


    keyword=Re: Start/End Run Script Status for d2n

    I have updated the EPICS loggers for the left HRS DAQ and tested them out.
    Everything appears to be working fine.

    I have also fixed the end-of-run scaler summary information for the left
    HRS and updated the calculation of the deadtime, i.e. I removed the
    correlations with T5 and T6. The deadtime now matches datamon reasonably

    A copy of this log entry has been emailed to: brads,moffit,rom